OpenStudy (anonymous):

What is the solution set for |–4x – 8| = 4? x = 3 and x = 1 x = 3 and x = –1 x = –3 and x = 1 x = –3 and x = –1

OpenStudy (anonymous):

I think its C

OpenStudy (anonymous):

Ok let us see..

OpenStudy (jasmineflvs):

yaw two or three opinions are needed

OpenStudy (anonymous):

|a| = b a = b or a = -b

OpenStudy (jasmineflvs):

explain it not give the answer yet

OpenStudy (anonymous):

It's a method, not the answer...

OpenStudy (anonymous):

\[ |–4x – 8| = 4 \rightarrow –4x – 8 = \pm4 \rightarrow –4x = \pm4 +8\] –4x = 4 +8 or –4x = -4 +8 -4x= 12 or -4x=4 x=12/(-4) or x= 4/(-4) x=3 or x=-1 Thus second option appears to be true

OpenStudy (jasmineflvs):

yup i thin so too

OpenStudy (anonymous):

Okay well im not sure if this is the right way to do it but -4x-8=4 +8 +8 -4x=12 divide by 4 divide by 4 x=3 0r 1

OpenStudy (anonymous):

yess i feel stupid -_- lol

zepdrix (zepdrix):

Hmm these are the answers im coming up with :[ \[\large -4x-8=4 \qquad\to\qquad x=-3\]\[\large -4x-8=-4 \qquad\to\qquad x=-1\]

zepdrix (zepdrix):

In your last step there, diving by 4 gives you\[-4x=12 \qquad\to\qquad -x=3\]

zepdrix (zepdrix):

dividing*

OpenStudy (anonymous):

So how do i solve these correctly how @dpasingh did it ?

zepdrix (zepdrix):

He did it incorrectly I'm afraid. :[\[\Large -4x-8=4\]Adding 8 to each side,\[\Large -4x=12\]dividing each side by -4,\[\Large x=-3\] The other solution,\[\Large -4x-8=-4\]Adding 8 to each side,\[\Large -4x=4\]Dividing each side by -4,\[\Large x=-1\]

OpenStudy (anonymous):

so i was kinda right ? lol

zepdrix (zepdrix):

With your C guess? :) lol Yah I guess so, you had one of the solutions correct at least ^^ hehe

zepdrix (zepdrix):

Err your last guess, where you showed the work ~ looks good, just don't forget about those negatives when you divide :)

OpenStudy (anonymous):

Okay lol thank you so much :)